Find all real solutions to the equation:
(x-4)^2=16

Answer by htmentor(1343)

Find all real solutions to the equation:
(x-4)^2=16
===============
x^2 - 8x + 16 - 16 = 0
x^2 - 8x = 0
Factor:
x(x-8) = 0
Solutions: x = 0, x = 8